hi - trying to help my middle schooler with this problem - i need help!
Flying with the wind, a pilot travels 600 miles twn 2 cities in 4 hours. The return trip into the wind takes 5 hours. The speed of the wind remains constant during the trip. Find the average speed of the plane with no wind and the speed of the wind.
Thank you!!

Answer by lwsshak3(11628)

Flying with the wind, a pilot travels 600 miles twn 2 cities in 4 hours. The return trip into the wind takes 5 hours. The speed of the wind remains constant during the trip. Find the average speed of the plane with no wind and the speed of the wind.
***
let x=speed of plane with no wind
let c=speed of wind
x+c=speed of plane with wind
x-c=speed of plane against wind
travel time*speed=distance
..
4(x+c)=600
5(x-c)=600
..
4x+4c=600
5x-5c=600
..
20x+20c=3000 (mult. by 5)
20x-20c=2400 (mult. by 4)
add
40x=5400
x=135
4c=600-4x=60
c=15
..
speed of plane with no wind=135 mph
speed of wind=15 mph